Jar A contains 5 red balls and 7 white balls. Jar B contains 8 red balls and 4 white balls. A fair die is rolled. If a 1 or 2 comes up, a ball is randomly selected from jar A, otherwise, a ball is randomly selected from jar B
a)Find the probability that a red ball is selected
b)Given that the ball selected is red, find the probability that it came from jar B

Answer by sudhanshu_kmr(1152)


a) 2/6 * 5/12 + 4/6 * 8/12
= 5/36 + 4/9
= 21/36
= 7/12

b) 4/6*8/12 / [ 2/6 * 5/12 + 4/6 * 8/12] ( by Bye's method)
= 4/9 / [7/12]
= 16/21